# Direct Imaging of Hydrogen-Driven Dislocation and Strain Field Evolution in a Stainless Steel Grain

### v2

@@ Acknowledgements
Conceptualization: D.Y., F.H.; Data Curation: D.Y., M.S.; Formal Analysis: D.Y., M.S., E.T., I.K.R., F.H.; Funding Acquisition: E.T., I.K.R., F.H.; Investigation: D.Y., G.H., K.S., R.J.H., W.C., D.N.; Methodology: D.Y., B.D., K.S., R.J.H., W.C.; Project Administration: F.H.; Resources: D.Y., G.H., K.S., A.M, F.H.; Software: D.Y., M.S., E.T., F.H.; Supervision: E.T., I.K.R., F.H.; Validation: D.Y., M.S.; Visualisation: D.Y., M.S.; Writing – Original Draft Preparation: D.Y., M.S., F.H.; Writing – Review & Editing: All authors.
@@
The diffraction patterns, reconstructions, and data analysis scripts for producing the figures will be publicly available at Zenodo: http://doi.org/10.5281/zenodo.14503567 after publication.

### v4

@@ Acknowledgements
D.Y. and F.H. conceptualized the study; D.Y. and M.S. curated the data; D.Y., M.S., E.T., I.K.R., and F.H. performed formal analysis; E.T., I.K.R., and F.H. acquired funding; D.Y., G.H., K.S., R.J.H., W.C., and D.N. performed investigation; D.Y., B.D., K.S., R.J.H., and W.C. were associated with methodology; F.H. was associated with project administration; D.Y., G.H., K.S., A.M, and F.H. acquired resources; D.Y., M.S., E.T., and F.H. were associated with the software; E.T., I.K.R., and F.H. supervised the study; D.Y. and M.S. performed validation; D.Y. and M.S. visualized the study; D.Y., M.S., and F.H. wrote the original draft; All authors reviewed and edited the final manuscript.
@@
The data that support the findings of this study are openly available in Zenodo [97].